OpenStudy (anonymous):

find all zeros. f(x)=(2x-1)(x-5)

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

x=-5 will make the second term zero (and thus the function zero). The first term we set equal to zero and solve: 2x-1=0 thus, x=1/2 will make the function zero as well. The solutions are x=1/2, x=-5

OpenStudy (anonymous):

oooops meant x=5 not -5

OpenStudy (anonymous):

could you make each step?

OpenStudy (anonymous):

ah I see

OpenStudy (anonymous):

so you use 2x-1=0 and x-5= 0 right?